Note the following (charges@ beginning of each month

1. Bought a car for 16250 and made down payment of 7.5%

2. Paid with visa for winter dress for 927

3. TTC bus pass for 6 months paid 311

4. Lawn mover for 1156

Had a meeting with Scotiabank to consolidate the debt into two classes

Class A the car loan plus bus pass

Class B the others.

If I want to pay class A over 6 years how much should I pay month at interest of 2% monthly


If am paying of class B by regular payments of 165 monthly how long will it take me to knock off the loan at 18% per annum.
1
Expert's answer
2022-02-01T14:38:34-0500

"A=16250-16250\times0.075+311=15342.25"

B=927+1156=2083


pay month class A:

"P=S(i+\frac{i}{(1+i)^n-1})"

"P=15342.25(0.02+\frac{0.02}{(1+0.02)^{12}-1})=1450.76"


paying of class B;


S=2083

P=150


"n=\frac{log\frac{S}{P}}{log(1+i)}"



"n=\frac{log\frac{2083}{165}}{log(1+0.18)}=\frac{log12.624242}{log1.18}=\frac{1.101}{ 0.07188}=15.3"